# MTLSO: A Multi-Task Learning Approach for Logic Synthesis Optimization

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ASPDAC ’25: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanisbn: 978-1-4503-XXXX-X/18/06
@@ 2.1. Logic Synthesis Optimization
Given the extreme complexity of VLSI chip designs, there has been a recent trend toward employing machine learning techniques to expedite design closure (Hamolia and Melnyk, 2021). In this context, Yu (Yu, 2023) examines the potential of machine learning models to enhance efficiency across various stages of EDA, including LSO. There are studies such as (Hosny et al., 2020) that compute representations of netlists by employing traditional techniques to extract hand-crafted features from AIGs. In contrast, other approaches utilize Graph Neural Networks (GNNs) as more advanced graph representation learners. These methodologies typically begin by employing a simple plain GNN to encode the input AIG. They subsequently employ various techniques to learn the representations of synthesis recipes. Prediction of QoR values for pairs of AIGs and recipes is then conducted by leveraging both the computed graph representations and the recipe representations. For instance, Chowdhury et al. (Chowdhury et al., 2021) learn representations of synthesis recipes by passing them through a set of 1D convolution layers. They then concatenate these representations with those learned for circuits by GNNs to predict QoR values. LOSTIN (Wu et al., 2022a) and GNN-H (Wu et al., 2022b) use LSTM to learn representations of synthesis recipes, capturing the relative ordering of logic transformations within them. Yang et al. (Yang et al., 2022a) adopt a similar strategy but replace the LSTM with the self-attention mechanism of the Transformer to learn representations of the recipes. None of these methods account for the varying importance of different graph nodes in learning the final graph-level representation of each AIG. They treat all graph nodes with equal importance, resulting in inefficiencies and reduced expressiveness of the final representation, particularly for very large AIGs, which are common in the LSO problem. Additionally, they have not addressed the critical issue of overfitting caused by severe data scarcity, which is a significant challenge in ML-based LSO solvers.
@@ 3.1. Problem Formulation
(1) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f:\mathcal{G}\times\mathcal{R}\rightarrow\mathbb{R}}$
@@ 3.2. Graph Encoder
(2) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{l+1}=\text{GNN}(X^{l},A^{l})}$
@@
(3) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{l+1},X^{l+1}=\text{{GraphDownsample}}(A^{l},H^{l+1})}$
where $A^{l+1}\in\{0,1\}^{N^{l+1}\times N^{l+1}}$ is the pruned adjacency matrix at the end of the $l$-th layer, and $X^{l+1}\in\mathbb{R}^{N^{l+1}\times F^{l+1}}$ is the representation matrix for the $N^{l+1}$ remaining nodes. These two matrices are then given as inputs to the next encoding layer. After passing $A^{0}$ and $X^{0}$ through $L$ consecutive graph encoding blocks, where the output of each block serves as the input for the subsequent block, the final outputs $A^{L}\in\{0,1\}^{N^{L}\times N^{L}}$ and $X^{L}\in\mathbb{R}^{N^{L}\times F^{L}}$ are computed. Here, $N^{L}$ represents the number of nodes remaining after all pruning iterations, while $F^{L}$ signifies the size of the final node representations. The final representation matrix $X^{L}$ is subsequently fed into a global pooling module to compute the graph. This process aggregates the representations learned for the most critical nodes after multiple steps of subsequent graph neural network operations and graph downsampling:
(4) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h_{G}=\text{{GraphPool}}(X^{L})}$
@@ 3.3.1. Embedding
(6) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{e}_{i}=\textsc{RecipeEmbed}(r_{i})}$
@@ 3.3.2. Convolutional Layers
(7) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda^{m}_{i}=\textsc{ConvLayer}^{m}(\mathbf{e}_{i})}$
@@
(8) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{i}=\text{Concat}(\lambda_{i}^{1},\lambda_{i}^{2},\ldots,\lambda_{i}^{M})}$
@@ 3.4.1. Binary Multi-label Graph Classification
(9) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}P^{G}_{\text{classification}}=\textsc{GraphClassify}(h_{G})}$
@@
(10) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{\text{classification}}=\textsc{BinaryCrossEntropy}(P^{G}_{\text{classification}},C^{G})}$
@@ 3.4.2. QoR Value Regression
(11) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y^{G}_{i}=\textsc{Decoder}\big{(}\text{Concat}(h_{G},\lambda_{i},P_{\text{classification}}^{G})\big{)}}$
@@
(12) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{\text{regression}}=\textsc{RegressionLoss}(y^{G}_{i},q^{G}_{i})}$
@@
(13) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}=\mathcal{L}_{\text{classification}}+\gamma\mathcal{L}_{\text{regression}}}$
@@ 4.5. Results
The main results, reported in Table 2, are averaged over multiple experimental runs. These findings demonstrate that our proposed MTLSO method outperforms the four baseline models in nearly all scenarios, with average gains of 8.22% in delay and 5.95% in area across all baselines and datasets. This underscores the critical advantage of employing a multi-task learning approach combined with a hierarchical graph encoder, rather than relying on simpler alternatives, for effectively addressing this problem.
@@
Delay Area
OpenABC-D EPFL CD OpenABC-D EPFL CD
Chowdhury et al. (Chowdhury et al., 2021) ($Y_{1}$) 23.66 $\pm$ 0.19 4.18 $\pm$ 0.07 15.88 $\pm$ 0.41 2.71 $\pm$ 0.02 2.46 $\pm$ 0.03 3.57 $\pm$ 0.10
LOSTIN (Wu et al., 2022a) ($Y_{2}$) 24.61 $\pm$ 0.03 3.96 $\pm$ 0.02 16.76 $\pm$ 0.13 2.35 $\pm$ 0.07 2.30 $\pm$ 0.01 3.46 $\pm$ 0.17
GNN-H (Wu et al., 2022b) ($Y_{3}$) 24.31 $\pm$ 0.27 3.96 $\pm$ 0.03 16.80 $\pm$ 0.06 2.33 $\pm$ 0.06 2.34 $\pm$ 0.02 3.59 $\pm$ 0.12
Yang et al. (Yang et al., 2022a) ($Y_{4}$) 24.58 $\pm$ 0.13 3.96 $\pm$ 0.02 17.09 $\pm$ 0.08 3.77 $\pm$ 0.00 2.39 $\pm$ 0.00 3.81 $\pm$ 0.05
MTLSO ($X$) 22.93 $\pm$ 0.23 3.94 $\pm$ 0.01 13.75 $\pm$ 0.25 2.57 $\pm$ 0.03 2.23 $\pm$ 0.04 3.33 $\pm$ 0.08
$\frac{Y_{1}-X}{Y_{1}}\times 100$ 3.09 5.74 13.41 5.17 9.35 6.72
Gain $\frac{Y_{2}-X}{Y_{2}}\times 100$ 6.83 0.51 17.96 - 9.36 3.04 3.76
$\frac{Y_{3}-X}{Y_{3}}\times 100$ 5.68 0.51 18.15 - 10.30 4.70 7.24
$\frac{Y_{4}-X}{Y_{4}}\times 100$ 6.71 0.51 19.54 31.83 6.69 12.60
It is important to note that our main architecture shares similarities with the one proposed by Chowdhury et al., particularly in our decision to use less complex techniques such as GCN (Kipf and Welling, 2016) as the core of our graph encoder and 1D convolution layers for encoding the recipes. This contrasts with the more advanced recipe encoders, like LSTM or Transformer, employed by LOSTIN (Wu et al., 2022a), GNN-H (Wu et al., 2022b), and Yang et al. (Yang et al., 2022a), which consider the ordering of logic transformations within each recipe. We opted for simpler model components to demonstrate that multi-task learning powered by HGRL, even with these basic components, can outperform the baselines. This suggests that performance could be further enhanced by using more advanced GNNs like GIN, as utilized by LOSTIN (Wu et al., 2022a), or by incorporating more sophisticated recipe encoders.
@@ 4.6. Ablation Study
Table 3. Ablation Study on Model Components.
@@
We performed an additional ablation study on the number of Graph Encoding/Downsampling layers, denoted as $L$. In addition to our main experiments with $L=2$, we investigated the cases where $L$ was set to $1$ and $3$. The results for delay minimization are presented in Figure 2. We assessed performance using graph classification metrics and MAPE to understand how the hyperparameter $L$ influences both the quality of the learned graph-level representation and the efficiency of the regression task. The results indicate that performance consistently improves with more than one encoding layer. This underscores the necessity of adopting a hierarchical strategy for effectively learning representations of large AIGs. Moreover, as the results suggest, the optimal $L$ is 2 for EPFL and CD, and 3 for OpenABC-D based on the graph classification metrics. This implies that we can further improve the quality of graph representations by using the optimal value of the hyperparameter $L$ for each individual dataset.
Figure 2. Ablation Study on the Number of Successive Graph Encoding/Downsampling Layers ($L$).
Finally, an ablation study was conducted on the ratio hyperparameter $\alpha$, which determines the percentage of nodes retained by the graph downsampling module during each graph encoding block. The results, illustrated in charts within Figure 3, encompass three values of $\alpha$: specifically, 0.5 alongside the extremes of 0.1 and 0.9. Across these experiments, optimal performance consistently favored $\alpha=0.5$, with both extreme values showing inferior results. This finding supports the importance of adopting such a hierarchical strategy for encoding AIGs, since a portion of nodes are less informative and should be filtered out (as evidenced by $\alpha=0.9$ performing worse than $\alpha=0.5$). Conversely, the underperformance of $\alpha=0.1$ highlights that some nodes have a direct influence on the quality of the final graph-level representation, and hence should not be pruned. This underscores the need to set an optimal value for this hyperparameter.
Figure 3. Ablation Study on the Node Retainment Ratio ($\alpha$) of the Graph Downsampling Module.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanbooktitle: 30th Asia and South Pacific Design Automation Conference (ASPDAC ’25), January 20–23, 2025, Tokyo, Japandoi: 10.1145/3658617.3697721isbn: 979-8-4007-0635-6/25/01ccs: Hardware Electronic design automation (EDA); Logic synthesisccs: Computing methodologies Machine learning Machine learning approaches
@@ 2.1. Logic Synthesis Optimization
Given the extreme complexity of VLSI chip designs, there has been a recent trend toward employing machine learning techniques to expedite design closure (Hamolia and Melnyk, 2021). In this context, Yu (Yu, 2023) examines the potential of machine learning models to enhance efficiency across various stages of EDA, including LSO. FlowTune (Neto et al., 2022) leverages a domain-specific, multi-stage multi-armed bandit approach to explore and optimize synthesis toolflows. There are studies such as (Hosny et al., 2020) that compute representations of netlists by employing traditional techniques to extract hand-crafted features from AIGs. In contrast, other approaches utilize Graph Neural Networks (GNNs) as more advanced graph representation learners. These methodologies typically begin by employing a simple plain GNN to encode the input AIG. They subsequently employ various techniques to learn the representations of synthesis recipes. Prediction of QoR values for pairs of AIGs and recipes is then conducted by leveraging both the computed graph representations and the recipe representations. For instance, Chowdhury et al. (Chowdhury et al., 2021) learn representations of synthesis recipes by passing them through a set of 1D convolution layers. They then concatenate these representations with those learned for circuits by GNNs to predict QoR values. LOSTIN (Wu et al., 2022a) and GNN-H (Wu et al., 2022b) use LSTM to learn representations of synthesis recipes, capturing the relative ordering of logic transformations within them. Yang et al. (Yang et al., 2022a) adopt a similar strategy but replace the LSTM with the self-attention mechanism of the Transformer to learn representations of the recipes. None of these methods account for the varying importance of different graph nodes in learning the final graph-level representation of each AIG. They treat all graph nodes with equal importance, resulting in inefficiencies and reduced expressiveness of the final representation, particularly for very large AIGs, which are common in the LSO problem. Additionally, they have not addressed the critical issue of overfitting caused by severe data scarcity, which is a significant challenge in ML-based LSO solvers.
@@ 3. Methodology
[…]
Overview of our MTLSO model which consists of Graph Encoder, Recipe Encoder, Binary Multi-label Graph Classifier, and Decoder. The model jointly optimizes graph classification and regression tasks.
@@ 3.1. Problem Formulation
(1) ${\color[rgb]{0,0,0}f:\mathcal{G}\times\mathcal{R}\rightarrow\mathbb{R}}$
@@ 3.2. Graph Encoder
(2) $H^{l+1}=\text{GNN}(X^{l},A^{l})$
@@
(3) $A^{l+1},X^{l+1}=\text{{GraphDownsample}}(A^{l},H^{l+1})$
where $A^{l+1}\in\{0,1\}^{N^{l+1}\times N^{l+1}}$ is the pruned adjacency matrix at the end of the $l$-th layer, and $X^{l+1}\in\mathbb{R}^{N^{l+1}\times F^{l+1}}$ is the representation matrix for the $N^{l+1}$ remaining nodes. These two matrices are then given as inputs to the next encoding layer. After passing $A^{0}$ and $X^{0}$ through $L$ consecutive graph encoding blocks, where the output of each block serves as the input for the subsequent block, the final outputs $A^{L}\in\{0,1\}^{N^{L}\times N^{L}}$ and $X^{L}\in\mathbb{R}^{N^{L}\times F^{L}}$ are computed. Here, $N^{L}$ represents the number of nodes remaining after all pruning iterations, while $F^{L}$ signifies the size of the final node representations. The final representation matrix $X^{L}$ is subsequently fed into a global pooling module to compute the graph representation. This process aggregates the representations learned for the most critical nodes after multiple steps of subsequent graph neural network operations and graph downsampling:
(4) $h_{G}=\text{{GraphPool}}(X^{L})$
@@ 3.3. Recipe Encoder
In this paper, we adopted the same set of transformations as described in (Chowdhury et al., 2021), where $m=7$ and the set $\{C_{1},C_{2},\ldots,C_{7}\}$ corresponds to Balance (b), Rewrite (rw, rw -z), and Refactor (rf, rf -z).
@@ 3.3.1. Embedding
(6) $\mathbf{e}_{i}=\textsc{RecipeEmbed}(r_{i})$
@@ 3.3.2. Convolutional Layers
(7) $\lambda^{m}_{i}=\textsc{ConvLayer}^{m}(\mathbf{e}_{i})$
@@
(8) $\lambda_{i}=\text{Concat}(\lambda_{i}^{1},\lambda_{i}^{2},\ldots,\lambda_{i}^{M})$
@@ 3.4.1. Binary Multi-label Graph Classification
(9) ${\color[rgb]{0,0,0}P^{G}_{\text{classification}}=\textsc{GraphClassify}(h_{G})}$
@@
(10) ${\color[rgb]{0,0,0}\mathcal{L}_{\text{classification}}=\textsc{BinaryCrossEntropy}(P^{G}_{\text{classification}},C^{G})}$
@@ 3.4.2. QoR Value Regression
(11) ${\color[rgb]{0,0,0}y^{G}_{i}=\textsc{Decoder}\big{(}\text{Concat}(h_{G},\lambda_{i},P_{\text{classification}}^{G})\big{)}}$
@@
(12) ${\color[rgb]{0,0,0}\mathcal{L}_{\text{regression}}=\textsc{RegressionLoss}(y^{G}_{i},q^{G}_{i})}$
@@
(13) ${\color[rgb]{0,0,0}\mathcal{L}=\mathcal{L}_{\text{classification}}+\gamma\mathcal{L}_{\text{regression}}}$
@@ 4.4. Implementation Details
In our experimental setup, two-thirds of the graphs in each dataset are randomly allocated for model training, while the remaining one-third is reserved for testing. The evaluation results, presented in the following subsections (i.e., Subsections 4.5 and 4.6), are reported on the test set.
@@ 4.5. Results
Metric Dataset Method Gain (%)
Chowdhury et al. (Chowdhury et al., 2021) ($Y_{1}$) LOSTIN (Wu et al., 2022a) ($Y_{2}$) GNN-H (Wu et al., 2022b) ($Y_{3}$) Yang et al. (Yang et al., 2022a) ($Y_{4}$) MTLSO ($X$) $\frac{Y_{1}-X}{Y_{1}}\times 100$ $\frac{Y_{2}-X}{Y_{2}}\times 100$ $\frac{Y_{3}-X}{Y_{3}}\times 100$ $\frac{Y_{4}-X}{Y_{4}}\times 100$
Delay OpenABC-D 23.66 $\pm$ 0.19 24.61 $\pm$ 0.03 24.31 $\pm$ 0.27 24.58 $\pm$ 0.13 22.93 $\pm$ 0.23 3.09 6.83 5.68 6.71
EPFL 4.18 $\pm$ 0.07 3.96 $\pm$ 0.02 3.96 $\pm$ 0.03 3.96 $\pm$ 0.02 3.94 $\pm$ 0.01 5.74 0.51 0.51 0.51
CD 15.88 $\pm$ 0.41 16.76 $\pm$ 0.13 16.80 $\pm$ 0.06 17.09 $\pm$ 0.08 13.75 $\pm$ 0.25 13.41 17.96 18.15 19.54
Area OpenABC-D 2.71 $\pm$ 0.02 2.35 $\pm$ 0.07 2.33 $\pm$ 0.06 3.77 $\pm$ 0.00 2.57 $\pm$ 0.03 5.17 -9.36 -10.30 31.83
EPFL 2.46 $\pm$ 0.03 2.30 $\pm$ 0.01 2.34 $\pm$ 0.02 2.39 $\pm$ 0.00 2.23 $\pm$ 0.04 9.35 3.04 4.70 6.69
CD 3.57 $\pm$ 0.10 3.46 $\pm$ 0.17 3.59 $\pm$ 0.12 3.81 $\pm$ 0.05 3.33 $\pm$ 0.08 6.72 3.76 7.24 12.60
The main results, reported in Table 2, are averaged over multiple experimental runs. These findings demonstrate that our proposed MTLSO method outperforms the four baseline models in nearly all scenarios, with average gains of 8.22% in delay and 5.95% in area across all baselines and datasets. This underscores the critical advantage of employing a multi-task learning approach combined with a hierarchical graph encoder, rather than relying on simpler alternatives, for effectively addressing this problem. More specifically, the integration of an additional classification task contributed significantly to these gains by introducing increased supervision during training. This strategy effectively leverages the limited yet highly valuable labeled data, thereby reducing overfitting and enhancing the model’s ability to generalize to unseen test data, as shown in Table 2.
It is important to note that our main architecture shares similarities with the one proposed by Chowdhury et al. (Chowdhury et al., 2021), particularly in our decision to use less complex techniques such as GCN (Kipf and Welling, 2016) as the core of our graph encoder and 1D convolution layers for encoding the recipes. This contrasts with the more advanced recipe encoders, like LSTM or Transformer, employed by LOSTIN (Wu et al., 2022a), GNN-H (Wu et al., 2022b), and Yang et al. (Yang et al., 2022a), which consider the ordering of logic transformations within each recipe. We opted for simpler model components to demonstrate that multi-task learning powered by HGRL, even with these basic components, can outperform the baselines. This suggests that performance could be further enhanced by using more advanced GNNs like GIN, as utilized by LOSTIN (Wu et al., 2022a), or by incorporating more sophisticated recipe encoders.
@@ 4.6. Ablation Study
Table 3. Ablation Study Results of Model Components in Terms of MAPE.
@@
We performed an additional ablation study on the number of Graph Encoding/Downsampling layers, denoted as $L$. In addition to our main experiments with $L=2$, we investigated the cases where $L$ was set to $1$ and $3$, keeping all other hyperparameters consistent with those reported in Subsection 4.4. The results for delay minimization are presented in Figure 2. We assessed performance using graph classification metrics and MAPE to understand how the hyperparameter $L$ influences both the quality of the learned graph-level representation and the efficiency of the regression task. The results indicate that performance consistently improves with more than one encoding layer. This underscores the necessity of adopting a hierarchical strategy for effectively learning representations of large AIGs. Moreover, as the results suggest, the optimal $L$ is 2 for EPFL and CD, and 3 for OpenABC-D based on the graph classification metrics. This implies that we can further improve the quality of graph representations by using the optimal value of the hyperparameter $L$ for each individual dataset.
[…]
Ablation Study results showing the effect of varying the number of successive Graph Encoding/Downsampling Layers (L) on the performance metrics. Each row of subplots corresponds to one dataset, and each column corresponds to a specific evaluation metric.
Figure 2. Ablation Study on the Number of Successive Graph Encoding/Downsampling Layers ($L$). The results are reported as percentages. Each row of subplots corresponds to one dataset, and each column corresponds to an evaluation metric.
Finally, an ablation study was conducted on the ratio hyperparameter $\alpha$, which determines the percentage of nodes retained by the graph downsampling module during each graph encoding block. The results, illustrated in charts within Figure 3, encompass three values of $\alpha$: specifically, 0.5 alongside the extremes of 0.1 and 0.9. All other hyperparameters were maintained as specified in Subsection 4.4 (e.g., $L=2$). Across these experiments, optimal performance consistently favored $\alpha=0.5$, with both extreme values showing inferior results. This finding supports the importance of adopting such a hierarchical strategy for encoding AIGs, since a portion of nodes are less informative and should be filtered out (as evidenced by $\alpha=0.9$ performing worse than $\alpha=0.5$). Conversely, the underperformance of $\alpha=0.1$ highlights that some nodes have a direct influence on the quality of the final graph-level representation, and hence should not be pruned. This underscores the need to set an optimal value for this hyperparameter.
[…]
Ablation Study on the Node Retainment Ratio (alpha) of the Graph Downsampling Module. Each row of subplots corresponds to one dataset, and each column corresponds to a specific evaluation metric.
Figure 3. Ablation Study on the Node Retainment Ratio ($\alpha$) of the Graph Downsampling Module. The results are reported as percentages. Each row of subplots corresponds to one dataset, and each column corresponds to an evaluation metric.